OpenStudy (anonymous):

solve for x: (4x-3)^2=-15

OpenStudy (anonymous):

so (4x-3)^2=(4x-3)(4x-3) right

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so take that and foil it

hero (hero):

(4x-3)(4x-3) = 4x(4x - 3) - 3(4x - 3) = 16x^2 - 12x - 12x +9 = 16x^2 - 24x + 9

OpenStudy (anonymous):

16x^2-24x+24=0

OpenStudy (anonymous):

but its not = to 0 its =-15

hero (hero):

(4x-3)(4x-3) = 4x(4x - 3) - 3(4x - 3) = 16x^2 - 12x - 12x +9 = 16x^2 - 24x + 9

OpenStudy (anonymous):

my bad see you already did that

OpenStudy (anonymous):

so now is there a common number that they are all divisible by?

OpenStudy (anonymous):

wait im at 16x^2-24x+9=-15 right?

OpenStudy (anonymous):

ya then add the 15 to get 16x^2-24x+24=0

OpenStudy (anonymous):

okay yeah i did that

OpenStudy (anonymous):

now find a common denominator

OpenStudy (anonymous):

8?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

so does that then turn into 2x^2-3x=3=0?

OpenStudy (anonymous):

2x^2-3x+3 sorry

hero (hero):

16x^2-24x+9=-15 16x^2 - 24x + 24 = 0 8(2x^2 - 3x + 3) = 0 2x^2 - 3x + 3 = 0 x^2 - 3/2 + 3/2 = 0 x^2 - 3/2x = -3/2 x^2 -3/2x + 9/16 = -3/2 + 9/16 (x - 3/4)^2 = -15/16 Solution is imaginary

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then what do i do?

OpenStudy (anonymous):

the answer is 3 + or - the square root of 15 divided by4 but idk how to get it

OpenStudy (anonymous):

i would do the quadratic formula

OpenStudy (anonymous):

-b + or - the square root of b^2 -4ac divided by 2a?

hero (hero):

\[\frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\]

OpenStudy (anonymous):

(-b+-sqrt(b^2-4ac))/2a

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

i got it thanks to both of you!

OpenStudy (anonymous):

not a problem

hero (hero):

You got what? Mind sharing what you got?

OpenStudy (anonymous):

3 + or - the square root of 15 divided by 4

OpenStudy (anonymous):

3+ or - i square root 15 divided by 4 sorry i forgot the "i"

OpenStudy (anonymous):

yup